OpenStudy (anonymous):

need help with math algerbra 2 its tha last question on the test i have been trying to solve it for hours i have been doing this 3 hours can not get this question commonnerfer 0 diana recived 65 points on a project for school she can make changes amd recive 3/10 of the missing points back she can make as many corrections as she wants create a formula for the sum of this geometric series and explain the steps in solving for the maximum grade

OpenStudy (anonymous):

What kind of math?

OpenStudy (anonymous):

diana recived 65 points on a project she can make changes ang get .3 of the missing points back she can fix it as many times as she wants. algerbra 2

OpenStudy (anonymous):

i think its a sigma notation

OpenStudy (solomonzelman):

Can you write the question clearly, the given and what you are asked to find.

OpenStudy (anonymous):

diana recived 65 points on a project for school she can make changes amd recive 3/10 of the missing points back she can make as many corrections as she wants create a formula for the sum of this geometric series and explain the steps in solving for the maximum grade

OpenStudy (anonymous):

that is the whole qusetion

OpenStudy (anonymous):

i assume it is out of 100

OpenStudy (anonymous):

She lost 35 points She can get 3/10 back 35*3/10=105/10=10.5 points maximum for new grade of 75.5 maximum. converging on 75.5 or 10.5 depending on your viewpoint

OpenStudy (anonymous):

i think its a sigma notation it can be redone infinitly

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

because u get the 75.5 but then u can redo it again and add a diffrent amount infintly